OpenStudy (anonymous):

Admission to the museum is $12 for children and $15 for adults. On a certain day, 2450 people go to the museum and $33,456 is collected. How many children and how many adults went to the museum? children adults

OpenStudy (anonymous):

@rydertheepic

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@wmj259

OpenStudy (anonymous):

2788 children

OpenStudy (wmj259):

Wow @rydertheepic

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Ops

OpenStudy (wmj259):

So they must have counted less children?

OpenStudy (wmj259):

@rydertheepic Okay you got your mistake

OpenStudy (wmj259):

33,456 = 12C+15A A=Adults C=Children

OpenStudy (anonymous):

Children 240

OpenStudy (anonymous):

CRAP

OpenStudy (anonymous):

I SUCK AT THIS

OpenStudy (anonymous):

stop giving me the straight answer ok

OpenStudy (anonymous):

i need toknow how to do this to understand it ok

OpenStudy (anonymous):

to know*

OpenStudy (igreen):

You can use this system of equations: 12x + 15y = 33456 x + y = 2450 Where x = children, y = adults

OpenStudy (igreen):

If you graph it you get a solution of (1098, 1352). So how many adults and children went? @RainbowBubbles01

OpenStudy (anonymous):

hold up

OpenStudy (wmj259):

This is an easy problem for a systems of equations Uploading picture

OpenStudy (anonymous):

you just gave me the answer i put 1098+1352 and gave me 2450

OpenStudy (anonymous):

right? or am i missing something?

OpenStudy (igreen):

No, 1098 and 1352 are your answers. (1098, 1352) (x, y) x = children y = adults Get it now? @RainbowBubbles01

OpenStudy (wmj259):

There you have it

OpenStudy (anonymous):

is that your boared?

OpenStudy (igreen):

x = children y = adults (1098, 1352) (x, y) 1098 = Children 1352 = Adults

OpenStudy (anonymous):

@wmj259 is that your bored? or boared how ever you spell it

OpenStudy (anonymous):

its true!!! :D

OpenStudy (wmj259):

Yes, my board where I make myself look like a nerd :(.

OpenStudy (anonymous):

put hi myawhith a heart and putit up then

OpenStudy (anonymous):

<3

OpenStudy (anonymous):

hi mya with a heart*

OpenStudy (wmj259):

Can't :'(

OpenStudy (anonymous):

then it not yours >:( LIER!!! xD

OpenStudy (igreen):

I like your picture..xD @wmj259

OpenStudy (wmj259):

@iGreen Thanks Your is really cute also :v) .

OpenStudy (wmj259):

thats a nose

OpenStudy (wmj259):

@RainbowBubbles01 my phones batter is below 5% gotta go pwoer-saving mode ;).

OpenStudy (igreen):

@wmj259 Thanks, I made it myself :P

OpenStudy (wmj259):

MEDAL!!!!!!!!

OpenStudy (igreen):

Lol